OpenStudy (anonymous):

The sequence is given by the following equation: t = (5 - n^2)/3n , where n E N [an element of a natural number]. Does this sequence have a limit of n approaches ∞?

OpenStudy (anonymous):

\[lim_{n\rightarrow\infty}\frac{5-n^2}{3n}\]

OpenStudy (anonymous):

if so, no limit because this thing goes to \[-\infty\]